OpenStudy (anonymous):

Given a generic equation for a parabola or an ellipse such as \(y=ax^2+bx+c\) or \(\frac{y^2}{a^2}+\frac{x^2}{b^2}=c^2\), how do you find the focus/foci?

OpenStudy (anonymous):

I guess the one for the ellipse should rather be just \(\frac{y^2}{a^2}+\frac{x^2}{b^2}=1\).

OpenStudy (vincent-lyon.fr):

For the ellipse, foci are: If a > b F1(c;0) and F2(-c;0) where c² = a² - b² If a < b F1(0;c) and F2(0;-c) where c² = b² - a²